OpenStudy (anonymous):

How many grams of ethane gas (C2H6) are in a 12.7 liter sample at 1.6 atmospheres and 24°C? Show all work used to solve this problem.

OpenStudy (anonymous):

Last question! Here's my answer: n = PV/RT P = 1.6 atm V = 12.7 L R = 0.0821 T = 24 degrees C ----> 297.15 degrees K n = (1.6 * 12.7) / (0.0821 * 297.15) n = 20.32 / 24.39 n = 0.83 mol C = 12.010 H = 1.0079 C2 = 12.010 * 2 = 24.02 H6 = 1.0079 * 6 = 6.0474 C2H6 = 30.0674 30.0674 * 0.83 = 24.95 grams (C2H6)

OpenStudy (anonymous):

@Cuanchi @ganeshie8

OpenStudy (anonymous):

@dan815 @paki

OpenStudy (cuanchi):

Good Job!!!

OpenStudy (anonymous):

Thank you! :)